OpenStudy (anonymous):

Analyze the cost purchase of two cars both four years old. One car costs Php 130,000 and gets 20 kilometers a gallon. The other costs Php 145,000 and gets 35 kilometers a gallon. Which car is the better buy if you plan to keep it for two years? Represent the answer graphically and algebraically.

OpenStudy (dumbcow):

Total cost = price of car + price of gas for 2 years Let k be total kilometers driven over 2 years Let g be price of 1 gallon of gas Car 1: Cost = 130,000 + kg/20 Car 2: Cost = 145,000 + kg/35 145,000 + kg/35 < 130,000 + kg/20 kg/35 - kg/20 < -15,000 4kg/140 - 7kg/140 < -15,000 -3kg < -2,100,000 kg > 700,000 k > 700,000/g Therefore Car2 is the better buy if you plan on driving more than (700,000/g) kilometers